OpenStudy (anonymous):

x over 2x-4-2 = 1 over x-2

OpenStudy (anonymous):

x/2x-6 = 1/ x-2 hint; cross multiply x(x-2)=1(2x-6) then solve

OpenStudy (anonymous):

its x over (2x-4) - 2=1 over x-2